Find the singleton sets and the pair of disjoint sets from the following



1. A = {x |x and x4 − 4 =0} and B = {y | y and x4 − 16 =0 =0}.

1
Expert's answer
2021-11-14T18:03:24-0500

Solution:

A = {x |x and x4 − 4 =0} and B = {y | y and y4 − 16 =0 =0}

"x^4 − 4 =0
\\ \Rightarrow x^4 =4
\\ \Rightarrow x^2=\pm2
\\ \Rightarrow x=\pm\sqrt{\pm2}
\\ \Rightarrow x=\sqrt{2},-\sqrt{2},\sqrt{2}i,-\sqrt{2}i"

And,

"y^4-16=0
\\\Rightarrow y^4=16
\\\Rightarrow y^2=\pm 4
\\\Rightarrow y=\pm \sqrt{\pm4}
\\\Rightarrow y=2, -2, 2i, -2i"

So, "A=\{\sqrt{2},-\sqrt{2},\sqrt{2}i,-\sqrt{2}i \}, B=\{2,-2,2i,-2i\}"

None is the singleton set (having just one element).

"\because A\cap B=\phi"

So, they are disjoint sets.